OpenStudy (anonymous):

Please, I need help, I really would like to understand what I am doing, cause then I can do it my self!!! :) Given: -2 times a number plus 20 is -114. Find the number. First write an equation that matches the statement above(4pts). Then, showing all your work, solve for the number (4pts). Make sure to show your check of your solution (2pts)

OpenStudy (perl):

We can label the unknown number x. This gives us the equation: -2*x + 20 = -114

OpenStudy (anonymous):

Let the number = x -2 times that number is = -2x

OpenStudy (anonymous):

plus 20 means add 20 to that number i.e. -2x + 20 and finally it says "is -114" or equals -114, so therefore you can rewrite it as: -2x + 20 = -114

OpenStudy (anonymous):

Ok, Perl I have tried doing this in reverse and whenever I get to adding the 20 it turns positive... would you know where I am going wrong?

OpenStudy (anonymous):

First step, get rid of - sign by multiplying across by -1 -1(-2x +20) = -1(-114)

OpenStudy (anonymous):

this gives you 2x -20 = 114

OpenStudy (anonymous):

Oh, hmm.... but tbh I am really confused :'(

OpenStudy (anonymous):

next bring across -20 to right hand side (and remember when you bring across change the sign) so 2x = 114 +20 2x = 134 divide both sides by 2 x = 67

OpenStudy (anonymous):

Oh....! See when I did it.. I would get -67 and then it would turn positive cause there was two negatives.

OpenStudy (anonymous):

well you want to get rid of the negative sign as early as possible, so that is why I multiplied across by -1 early on

OpenStudy (anonymous):

Oh ok.... That and this is making a lot of sense now!!! :)

OpenStudy (anonymous):

good, delighted to hear it!!

OpenStudy (anonymous):

the rule is try and get rid of negatives and fractions as early as possible

OpenStudy (anonymous):

Good to know!!! :) Could you possibly write it out for I can see it all in one thing? and is there away I can reward you for helping me?

OpenStudy (anonymous):

no probs

OpenStudy (anonymous):

Thanks :)

OpenStudy (anonymous):

minus 2 times (a number) = -2(x) plus 20 = -2(x) + 20 is (equals): -114 i.e -2(x) +20 = -114 -2x + 20 = -114 get rid of negative by x -1 as follows: -1(-2x + 20) = -1(-114) 2x -20(+20) = 114(+20) 2x = 114 + 20 2x = 134 2x/2 = 134/2 x = 67

OpenStudy (anonymous):

Thank you so much!!!! :)